OpenStudy (anonymous):

use an addition or subtraction formula to find the exact value of sin (-5pi/12)

OpenStudy (jdoe0001):

\(\bf -\cfrac{5}{12}\implies \begin{cases} \cfrac{\cancel{ 4 }}{\cancel{ 12 }}-\cfrac{\cancel{ 9 }}{\cancel{ 12 }} \\ \quad \\ \end{cases}\qquad thus \\ \quad \\ sin\left(-\cfrac{5\pi}{12}\right) \iff sin\left(\cfrac{\pi}{3}-\cfrac{3\pi}{4}\right)\)

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\bf -\cfrac{5}{12}\implies \begin{cases} \cfrac{\cancel{ 4 }}{\cancel{ 12 }}-\cfrac{\cancel{ 9 }}{\cancel{ 12 }} \\ \quad \\ \bf \cfrac{1}{3}-\cfrac{3}{4} \end{cases}\qquad thus \\ \quad \\ sin\left(-\cfrac{5\pi}{12}\right) \iff sin\left(\cfrac{\pi}{3}-\cfrac{3\pi}{4}\right)\)